OpenStudy (anonymous):

simplfiy. 2/x+1/9=1/3 . I know the answer is 9 but how?

OpenStudy (anonymous):

to clear things up its 2 over x plus 1 over 9

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

\[\frac{2}{x}=\frac{3}{9}-\frac{1}{9}=\frac{2}{9}\]

OpenStudy (anonymous):

2/x = 1/3 - 1/9 2/x = 2/9 Cross-multiply 18 = 2x x = 9

OpenStudy (anonymous):

what monebird said lol

OpenStudy (anonymous):

equation editor takes forever unless you type it manually =/